OpenStudy (anonymous):

find global max and min of f(x) = x-2ln (x-1),0

OpenStudy (anonymous):

sorry there must be some mistake here domain of \(\ln(x-1)\) is \(x>1\)

OpenStudy (anonymous):

how about \[f(x) x-2\ln(x-1), 0\le x \le2 \]

OpenStudy (anonymous):

the problem is that the domain of your function does not include \((0,1]\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

f(x)= x-2ln(x+1)\[f(x)= x-2\ln(x+1)0\le x \le2 \]

OpenStudy (anonymous):

help please!!